OpenStudy (anonymous):

lim x to 4 = sqrt{x+5} - sqrt5 / x-4

OpenStudy (anonymous):

what do you get when you plug in 4?

OpenStudy (anonymous):

looks like a direct substitution gives you \[\frac{\sqrt9-\sqrt5}{4-4}=\frac{3-\sqrt5}{0}\]

OpenStudy (anonymous):

since the denominator is zero and the numerator is not, there is not limit